A circle is divided into 9 uneven "pie" shaped pieces. Each section is labeled 1, 2, ... , 9. The circle is then spun and the number pointing north recorded. Estimate the central angle of the section labeled 4.
4 3 2 1 9 8 5 6 7 4 3 2 1 5 8 6 5 4

The probability of event A occurring is 0.3 and the probability of event B occurring is 0.4. What is P(A and B) given that A and B are independent events?

Of the 84 people who answered "yes" to a question, 14 were male. Of the 73 people that answered "no" to the question, 12 were male. If one person is selected at random from the group, what is the probability that the person answered "yes" or was male? Round to the nearest thousandth.
thank you!

Answer by ikleyn(52874)

.

            There is a RULE in this forum

            One and ONLY ONE problem/question per post.


            Please DO NOT violate this rule.

            It was established for your best interests.


~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~

I will answer one question only.

            The probability of event A occurring is 0.3 and the probability of event B occurring is 0.4.
            What is P(A and B) given that A and B are independent events?

Since the events  A  and  B  are independent,  


    P(A and B) = P(A)*P(B) = 0.3*0.4 = 0.12.


It is the definition of the property of two events to be independent (!)